40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6770     Entry time: Wed Jun 6 19:46:46 2012
Author: Suresh 
Type: Summary 
Category: IOO 
Subject: Tip-tilt assembly: current status and work remaining 

 

Recent History

The lower blades which I had given to the Physics Workshop for making a vacuum relief hole (using a sinker-EDM process) came back about ten days ago.   Merih Eken <meken@caltech.edu>,  the supervisor at the Physics Dept workshop, handled this matter for us.  The blades were sent to a local EDM machineshop and returned in about three working days ( a weekend intervened). 

IMG_0685.JPG  IMG_0687.JPG

Bob cleaned and handed them over to me yesterday evening.  

Current status

Today I have reassembled the four tip-tilts.  I have repacked them in clean bags (double bagged) shifted them to Clean Optics Cabinet (near the ETMX chamber).  The four tip-tilts are in the bottom-most shelf in the cabinet.  There are also some tip-tilt spares in a separate envelope.

Note:  The mirror holder is now held tightly by the eddy current dampers.  This was done for safety of the wires during transportation from LHO.  The eddy current damper in the front of the mirror has to be retracted to allow the mirror holder to swing free.  It has be to about 1mm away from the suspended mirror holder

Work Remaining

1) We need to install the quadrapus cables.  The connector placement on the BOSEM side will have some issues.  It is best to loosen the BOSEM seating as well as the connector seating screws and then push the cable connector into place.  Caution:  when the connector seating screws on the BOSEM are loosened the flexible ckt could be damaged by the loose connector.

2) Insert the mirrors into the mirror holders and balance the suspension such that the mirror's hang vertical and do not have a large yaw offset.

3) Adjust the wire suspension point height so that the flags are in the center of the BOSEM aperture.  Else they will strike against the

4) We need to adjust the position of the BOSEMs such that the shadow sensor signals are at 50%.  This ensures that all the magnets hang at an appropriate distance from their respective coils.

5) To do (3) we need to set up a shadow sensor read-out set-up for one tip-tilt (four sensors)

 

Attachment 2: IMG_0687.JPG  2.121 MB  | Hide | Hide all
IMG_0687.JPG
ELOG V3.1.3-